# Explainable AI for Classifying UTI Risk Groups Using a Real-World Linked EHR and Pathology Lab Dataset

### v3

@@
Abstract
The use of machine learning and AI on electronic health records (EHRs) holds substantial potential for clinical insight. However, this approach faces challenges due to data heterogeneity, sparsity, temporal misalignment, and limited labeled outcomes. In this context, we leverage a linked EHR dataset of approximately one million de-identified individuals from Bristol, North Somerset, and South Gloucestershire, UK, to characterize urinary tract infections (UTIs). We implemented a data pre-processing and curation pipeline that transforms the raw EHR data into a structured format suitable for developing predictive models focused on data fairness, accountability and transparency. Given the limited availability and biases of ground truth UTI outcomes, we introduce a UTI risk estimation framework informed by clinical expertise to estimate UTI risk across individual patient timelines. Pairwise XGBoost models are trained using this framework to differentiate UTI risk categories with explainable AI techniques applied to identify key predictors and support interpretability. Our findings reveal differences in clinical and demographic predictors across risk groups. While this study highlights the potential of AI-driven insights to support UTI clinical decision-making, further investigation of patient sub-strata and extensive validation are needed to ensure robustness and applicability in clinical practice.
@@ 5.1 Data Preparation
For the control group, consisting of individuals with a UTI likelihood of 0, we sampled individuals in a manner that aligned with the temporal distribution of the UTI group (likelihood $>$ 0) across the years. This ensures that both the UTI and control groups had similar distribution of observation time-frames, facilitating a fair comparison between the two groups with respect to the distribution of timelines across the years and seasons.
@@ 6.1 Descriptive Analysis
Control Group [^1] Individuals with a UTI likelihood $=$ 0 UTI Group [^2] Individuals with a UTI likelihood $>$ 0
@@
Figure 4: Age and sex distribution of the UTI group (likelihood $>$ 0). (Left) Age and sex distribution. (Right) Age distribution of different UTI likelihood groups.
@@ 6.2 Model Performance and Feature Importance
The SHAP summary plots (Figure 7) reveal evolving patterns in feature importance as the risk groups progress, reflecting both clinical and demographic factors that influence UTI likelihood. In the 0.0 vs. 0.2 model, features related to prior clinical interventions, such as the number of previous antibiotics dispensed and the number of urine specimens processed, play a prominent role in distinguishing between low-risk groups, suggesting that historical clinical events within one year contribute to predicting risk in these less certain cases. In the 0.4 vs. 0.6 model, the focus shifted towards more specific clinical features, including the number of UTI-specific antibiotics dispensed. Additionally, comorbidities like dementia and prior resistance discoveries become more relevant, indicating their growing importance in identifying patients at moderate risk. In the 0.2 vs. 0.6 model, which differentiates between low and moderately high-risk patients, the number of UTI-specific antibiotics dispensed again stands out as a key feature. For the 0.6 vs. 0.8 and 0.8 vs 1.0 models, age displays a clearer trend, with older patients more likely to be classified into the higher-risk groups. Features such as prior resistance discoveries and the number of urine specimens processed are particularly influential in the model of the 0.6 vs. 0.8. When distinguishing patients in the group of 0.8 vs. 1.0, the model highlights severe case predictors, such as housebound status and the number of previous hospital visits, underscoring the role of these factors in patients with hospital-confirmed UTI diagnoses (UTI likelihood $=$ 1.0).

### v4

@@
11institutetext: Yujie Dai 22institutetext: University of Bristol, UK 22email: yujie.dai@bristol.ac.uk 33institutetext: Brian Sullivan 44institutetext: University of Bristol, UK 44email: brian.sullivan@bristol.ac.uk 55institutetext: Axel Montout 66institutetext: University of Bristol, UK 66email: axel.montout@bristol.ac.uk 77institutetext: Amy Dillon 88institutetext: University of Bristol, UK 88email: amy.dillon@bristol.ac.uk 99institutetext: Chris Waller 1010institutetext: Kettering General Hospital NHS Foundation Trust, UK 1010email: christopher.waller@nhs.net 1111institutetext: Peter Acs 1212institutetext: NHS England, UK 1212email: peter.acs@nhs.net 1313institutetext: Rachel Denholm 1414institutetext: University of Bristol, UK 1414email: r.denholm@bristol.ac.uk 1515institutetext: Philip Williams 1616institutetext: University Hospitals Bristol and Weston NHS Foundation Trust, UK 1616email: Philip.Williams2@UHBW.nhs.uk 1717institutetext: Alastair D Hay 1818institutetext: University of Bristol, UK 1818email: alastair.hay@bristol.ac.uk 1919institutetext: Raul Santos-Rodriguez 2020institutetext: University of Bristol, UK 2020email: enrsr@bristol.ac.uk 2121institutetext: Andrew Dowsey 2222institutetext: University of Bristol, UK 2222email: andrew.dowsey@bristol.ac.uk
Explainable AI for Classifying UTI Risk Groups Using a Real-World Linked EHR and Pathology Lab Dataset
[…]
Axel Montout and
Amy Dillon and
Chris Waller and
Peter Acs and
Rachel Denholm and
Philip Williams and
Alastair D Hay and
[…]
@@ 5.1 Data Preparation
For the control group, consisting of individuals with a UTI likelihood of 0, we sampled individuals in a manner that aligned with the temporal distribution of the UTI group (likelihood >> 0) across the years. This ensures that both the UTI and control groups had similar distribution of observation time-frames, facilitating a fair comparison between the two groups with respect to the distribution of timelines across the years and seasons.
@@ 6.1 Descriptive Analysis
Control Group [^1] Individuals with a UTI likelihood == 0 UTI Group [^2] Individuals with a UTI likelihood >> 0
@@
Figure 4: Age and sex distribution of the UTI group (likelihood >> 0). (Left) Age and sex distribution. (Right) Age distribution of different UTI likelihood groups.
@@ 6.2 Model Performance and Feature Importance
The SHAP summary plots (Figure 7) reveal evolving patterns in feature importance as the risk groups progress, reflecting both clinical and demographic factors that influence UTI likelihood. In the 0.0 vs. 0.2 model, features related to prior clinical interventions, such as the number of previous antibiotics dispensed and the number of urine specimens processed, play a prominent role in distinguishing between low-risk groups, suggesting that historical clinical events within one year contribute to predicting risk in these less certain cases. In the 0.4 vs. 0.6 model, the focus shifted towards more specific clinical features, including the number of UTI-specific antibiotics dispensed. Additionally, comorbidities like dementia and prior resistance discoveries become more relevant, indicating their growing importance in identifying patients at moderate risk. In the 0.2 vs. 0.6 model, which differentiates between low and moderately high-risk patients, the number of UTI-specific antibiotics dispensed again stands out as a key feature. For the 0.6 vs. 0.8 and 0.8 vs 1.0 models, age displays a clearer trend, with older patients more likely to be classified into the higher-risk groups. Features such as prior resistance discoveries and the number of urine specimens processed are particularly influential in the model of the 0.6 vs. 0.8. When distinguishing patients in the group of 0.8 vs. 1.0, the model highlights severe case predictors, such as housebound status and the number of previous hospital visits, underscoring the role of these factors in patients with hospital-confirmed UTI diagnoses (UTI likelihood == 1.0).